OpenStudy (izzy18970):

How do you make 20/21 as a whole number?

OpenStudy (mertsj):

It is not a whole number so it cannot be written as a whole number.

OpenStudy (izzy18970):

ok

OpenStudy (anonymous):

you could say (1/21)*20, but you still have a rational part

OpenStudy (izzy18970):

it is just this is my problem.... The problem 4/3 / 3/5 has been put on the board. What multiplication problem can you write to find the quotient? What is the quotient of 4/7 / 3/5?

OpenStudy (izzy18970):



OpenStudy (zzr0ck3r):

\[\frac{\frac{a}{b}}{\frac{d}{k}}=\frac{a}{b}*\frac{k}{d}\]

OpenStudy (izzy18970):

ok

OpenStudy (izzy18970):

so you wouldn't change 3/5 into a reciprocal?

OpenStudy (mertsj):